### v2

# ARCE: Augmented RoBERTa with Contextualized Elucidations for NER in Automated Rule Checking

Abstract
Accurate information extraction from specialized texts is a critical challenge, particularly for named entity recognition (NER) in the architecture, engineering, and construction (AEC) domain to support automated rule checking (ARC). The performance of standard pre-trained models is often constrained by the domain gap, as they struggle to interpret the specialized terminology and complex relational contexts inherent in AEC texts. Although this issue can be mitigated by further pre-training on large, human-curated domain corpora, as exemplified by methods like ARCBERT, this approach is both labor-intensive and cost-prohibitive. Consequently, leveraging large language models (LLMs) for automated knowledge generation has emerged as a promising alternative. However, the optimal strategy for generating knowledge that can genuinely enhance smaller, efficient models remains an open question. To address this, we propose ARCE (augmented RoBERTa with contextualized elucidations), a novel approach that systematically explores and optimizes this generation process. ARCE employs an LLM to first generate a corpus of simple, direct explanations, which we term Cote, and then uses this corpus to incrementally pre-train a RoBERTa model prior to its fine-tuning on the downstream task. Our extensive experiments show that ARCE establishes a new state-of-the-art on a benchmark AEC dataset, achieving a Macro-F1 score of 77.20%. This result also reveals a key finding: simple, explanation-based knowledge proves surprisingly more effective than complex, role-based rationales for this task. The code is publicly available at: https://github.com/nxcc-lab/ARCE.
Index Terms—  Automated Rule Checking, Large Language Models, Named Entity Recognition.
1 INTRODUCTION
Fig. 1: The overall architecture of our proposed ARCE approach. The approach consists of three stages. Stage 1: We employ a LLM (e.g., Qwen3) to generate a Cote (contextualized task-oriented elucidation) corpus from raw domain texts using specialized prompts. Stage 2: A pre-trained RoBERTa-$w$-ext model undergoes incremental pre-training on the Cote corpus via a masked language modeling (MLM) objective, yielding the enhanced ARCE model. Stage 3: The ARCE model is augmented with a CRF layer and fine-tuned on the downstream NER task, and is then used for inference on new, unseen texts.
The architecture, engineering, and construction (AEC) industry, a cornerstone of the global economy, is laden with complex safety and regulatory requirements documented in vast volumes of unstructured text [1, 2, 3]. To navigate this complexity, automated rule checking (ARC) has emerged as a vital technology for ensuring regulatory compliance [4, 5]. A critical bottleneck in developing effective ARC systems, however, lies in the ”rule interpretation” stage, which necessitates the accurate extraction of semantic information from text—a task formally known as named entity recognition (NER) [6].
Early attempts to address this NER challenge involved fine-tuning pre-trained language models such as BERT [7] and RoBERTa [8] on small, in-domain labeled datasets. However, the performance of these models is often constrained by the significant domain gap between the general-purpose corpora used for their initial training and the specialized lexicon of the AEC domain [9, 10]. This gap makes robust and accurate information extraction a persistent challenge.
To bridge this domain gap, two primary paradigms have been explored. The first, exemplified by ARCBERT [2], involves further pre-training on large, human-curated, domain-specific corpora. While effective, this approach is exceptionally labor-intensive and costly. A more recent paradigm leverages large language models (LLMs) to automatically generate task-oriented knowledge, thereby augmenting smaller, more efficient models [11, 12]. Our work builds upon and significantly extends this latter approach.
We propose ARCE (augmented RoBERTa with contextualized elucidations), a novel approach that leverages an LLM (Qwen3-8B [13]) to generate a high-quality, task-oriented knowledge corpus—which we term Cote. This corpus is then used to incrementally pre-train a RoBERTa model, adapting it to the specialized domain before fine-tuning it on the downstream NER task. Our extensive experiments validate the superiority of the ARCE approach, establishing a new state-of-the-art on a benchmark AEC dataset. More significantly, our work reveals a counter-intuitive insight: the style of the generated knowledge is a critical factor, with simpler, direct explanations unexpectedly outperforming more complex, functional rationales. This finding distinguishes ARCE from other generative approaches like NuNER [12].
The main contributions of this paper are summarized as follows:
  • We propose ARCE, a novel approach that leverages LLM-generated elucidations to effectively adapt a RoBERTa model to a specialized domain, circumventing the need for large, human-curated corpora required by methods like ARCBERT.
  • We conduct a systematic study of knowledge generation strategies and discover that for this task, simple, explanation-based elucidations are significantly more effective than complex, role-based ones, offering a valuable ”less is more” insight for generative data augmentation.
  • We establish a new state-of-the-art performance on the benchmark dataset for NER in the ARC domain, outperforming a range of strong baselines, including domain-adapted models and fine-tuned LLMs.
2 METHODOLOGY
Our proposed approach, ARCE (augmented RoBERTa with contextualized elucidations), is designed to enhance NER performance in specialized domains by leveraging knowledge generated from a large language model, specifically Qwen3-8B [13]. As illustrated in Figure 1, the ARCE approach is decomposed into three sequential stages: (1) the generation of a Contextualized oriented task-elucidation (Cote) corpus; (2) incremental pre-training of a RoBERTa model on this corpus; and (3) downstream NER fine-tuning with a Conditional Random Field (CRF) layer.
Table 1: Results comparison with different methods. Bolded values indicate the best performance.
Models Macro-F1 with Strict Match (%) Macro-F1 with Partial Match (%)
Precision Recall Macro-F1 Precision Recall Macro-F1
Traditional Methods
NuNER-CRF [12] 64.61 62.79 63.69
RoBERTa-CRF [8] 66.94 62.35 64.56 /
BERT-base-Chinese-CRF [7] 77.54 71.02 72.93
RoBERTa-www-ext-CRF [14] 74.62 68.36 71.19
8B Large Language Models
Llama3.1-8B-Instruct [15] 21.71 10.65 13.00 31.93 20.31 22,91
Ministral-8B-Instruct 22.31 16.50 18.87 35.18 33.61 34.11
Qwen3-8B [13] 29.09 21.98 25.01 61.93 30.77 35.73
Large Language Models for Reasoning
DS-R1-Distill-Llama-8B [16] 13.79 9.63 10.51 34.90 23.56 25.39
DS-R1-Distill-Qwen-7B [16] 19.26 10.00 12.47 31.86 17.41 21.23
DS-R1-0528-Qwen3-8B [16] 25.98 11.37 15.60 42.11 15.73 21.99
Qwen3-8B-think [13] 41.28 28.77 32.55 55.16 36.73 41.68
Qwen3-235B-A22B-Instruct-2507 [17] 29.98 28.14 28.65 71.98 46.08 49.13
Qwen3-8B-SFT [17] 75.48 75.39 75.23 84.07 83.41 83.49
ARCBERT-CRF [2] 77.31 75.26 75.75 /
ARCE(Ours) 77.51 77.30 77.20 /
2.1 Stage 1: Contextualized Task-Oriented Elucidation (Cote) Generation
The cornerstone of our approach is the creation of a high-quality, task-specific corpus, which we term the Corpus of contextualized task-oriented elucidations ($\mathcal{C}_{Cote}$). Unlike traditional data augmentation methods that rely on generic external corpora, we employ a powerful LLM, denoted as $\mathcal{M}$, to generate elucidations explicitly tailored to the reasoning process of the NER task.
Given a sentence $x=\{w_{1},w_{2},\dots,w_{n}\}$ from a source dataset and an entity within it, defined by its text span $s$ and entity type $t$, we construct a specialized prompt $P$ based on a prompting strategy $\Pi$, such that $P=\Pi(x,s,t)$. This strategy directs the LLM to generate a textual elucidation $k$ that clarifies why the span $s$ corresponds to the type $t$ within the context of $x$, or describes the functional role it plays.
The generation of the elucidation text $k=\{k_{1},k_{2},\dots,k_{L}\}$ is an auto-regressive process, where the probability of the sequence is conditioned on the prompt $P$ and the LLM’s parameters $\theta_{\mathcal{M}}$. This is formally expressed as:
$p(k|P;\theta_{\mathcal{M}})=\prod_{i=1}^{L}p(k_{i}|k_{<i},P;\theta_{\mathcal{M}})$ (1)
By iterating this process over all entities in our source data, we compile the final Cote corpus, $\mathcal{C}_{Cote}$, which serves as the foundation for adapting our NER model:
$\mathcal{C}_{Cote}=\{k_{j}|k_{j}\sim p(k|\Pi(x_{j},s_{j},t_{j});\theta_{\mathcal{M}})\}$ (2)
where $(x_{j},s_{j},t_{j})$ represents the $j$-th entity instance in the source dataset.
2.2 Stage 2: Incremental Pre-training on Cote
To effectively integrate the domain-specific knowledge encapsulated in the elucidations into our NER model, we perform incremental pre-training on a RoBERTa model [14]. Let its initial parameters be $\theta_{RoBERTa}$. This stage adapts the general-domain model to the specific vocabulary, syntax, and reasoning patterns present in $\mathcal{C}_{Cote}$.
The training objective for this phase is the standard masked language modeling (MLM) task [7]. For each elucidation text $k\in\mathcal{C}_{Cote}$, we generate a corrupted version $\tilde{k}$ by randomly masking a subset of its tokens. Let $k_{masked}$ be the set of original tokens at the masked positions. The MLM loss, $\mathcal{L}_{MLM}$, is defined as the negative log-likelihood of correctly predicting these original tokens from the corrupted context $\tilde{k}$:
$\mathcal{L}_{MLM}(\theta)=-\sum_{k\in\mathcal{C}_{Cote}}\sum_{w\in k_{masked}}\log p(w|\tilde{k};\theta)$ (3)
The RoBERTa model’s parameters are updated by minimizing this loss function over the entire Cote corpus. This procedure yields an enhanced set of parameters, $\theta_{ARCE}$, which now encode both general linguistic understanding and the targeted task-specific knowledge, as shown in Eq. (4).
$\theta_{ARCE}=\arg\min_{\theta_{RoBERTa}}\mathcal{L}_{MLM}(\theta_{RoBERTa})$ (4)
2.3 Stage 3: Downstream NER Fine-tuning with CRF
The final stage involves fine-tuning the adapted ARCE model for the downstream NER task. To effectively model dependencies between adjacent labels, we augment the architecture with a conditional random field (CRF) layer. The CRF layer computes a score $s(x,y)$ for a label sequence $y$ by combining emission scores from the encoder and a learned transition matrix.
$s(x,y)=\sum_{i=1}^{n}E_{i,y_{i}}+\sum_{i=1}^{n-1}A_{y_{i},y_{i+1}}$ (5)
The model is jointly optimized by minimizing the negative log-likelihood of the ground-truth sequence $y_{\text{true}}$.
$p(y|x)=\frac{\exp(s(x,y))}{\sum_{y^{\prime}\in\mathcal{Y}(x)}\exp(s(x,y^{\prime}))}$ (6)
$\mathcal{L}_{\text{NER}}=-\log p(y_{\text{true}}|x)$ (7)
During inference, the Viterbi algorithm efficiently finds the label sequence $\hat{y}$ that maximizes this score.
$\hat{y}=\underset{y^{\prime}\in\mathcal{Y}(x)}{\arg\max}\,s(x,y^{\prime})$ (8)
3 EXPERIMENTS
3.1 Experimental Setup
We evaluate our approach on the public AEC-domain NER dataset by Zhou et al. [6], which contains 4336 entities across 611 sentences. Following standard protocol [2], we use an 8:1:1 train/validation/test split and report the Macro-F1 score as the primary metric to account for label imbalance.
In the incremental pre-training stage, we train the RoBERTa-www-ext model on our Cote corpus for 5 epochs with a learning rate of 5e-5, a weight decay of 0.01, and a batch size of 16. For the downstream NER fine-tuning, we train the ARCE model for 10 epochs using an AdamW optimizer and a linear learning rate scheduler. We employ a differential learning rate strategy, setting the learning rate for the CRF layer to 5e-1 and for the rest of the model to 5e-5. The batch size for fine-tuning is also set to 16.
3.2 Baselines
We benchmark ARCE against three categories of baseline models: (1) traditional fine-tuned encoders (e.g., BERT-CRF [7], RoBERTa-CRF [8]); (2) zero-shot LLMs (e.g., Qwen3-8B [13]); and (3) domain-specific SOTA models. Key SOTA baselines include Qwen3-8B-SFT [17], representing direct LLM fine-tuning; ARCBERT [2], which relies on a large human-curated corpus for pre-training; and NuNER [12], which employs an alternative LLM-based data generation strategy.
3.3 Main Results and Analysis
The primary experimental results are presented in Table I. Our approach, ARCE, establishes a new SOTA on the benchmark dataset, achieving a strict-match Macro-F1 score of 77.20%.
ARCE significantly outperforms traditional methods, surpassing BERT-base-Chinese-CRF by 4.27 points. In contrast, zero-shot LLMs perform poorly; for instance, Qwen3-8B scores only 25.01% under the strict-match criterion. Given that generative models often struggle with precise entity boundary detection [18, 19], we also evaluate them using a partial-match metric. This raises Qwen3-8B’s score to 35.73% and the fine-tuned Qwen3-8B-SFT’s score to 83.49%, yet the performance of zero-shot LLMs still highlights the critical need for domain-specific adaptation.
Crucially, under the strict-match criterion, ARCE also surpasses the most competitive SOTA baselines. It achieves a higher Macro-F1 score than both ARCBERT-CRF, which relies on a human-curated corpus, and the directly fine-tuned Qwen3-8B-SFT. This finding validates our core hypothesis: augmenting a smaller model with targeted, LLM-generated elucidations is a more effective and efficient strategy than relying on massive domain corpora or brute-force fine-tuning of larger models.
4 Discussion
Fig. 2: Results of the ablation experiment.
4.1 Ablation on Knowledge Generation Strategy
To isolate the impact of the knowledge generation strategy, we conduct an ablation study comparing our main model, ARCE, with two variants. These variants use Cote corpora generated from different prompts: ARCE-think, which produces verbose, step-by-step rationales, and ARCE-role, which analyzes an entity’s functional role. As shown in Figure 2, our standard, explanation-based ARCE achieves the highest Macro-F1 score, outperforming ARCE-think and ARCE-role. A deeper analysis reveals a precision-recall trade-off: ARCE-role obtains the highest precision (78.73%) at the cost of the lowest recall (72.88%), suggesting its complex rationales are overly specific and hinder generalization. In contrast, our approach strikes a superior balance between precision (77.51%) and recall (77.30%). This finding confirms that for domain adaptation in this context, a large volume of simple, direct explanations is more effective than fewer, more complex ones, robustly supporting a ”less is more” principle.
Fig. 3: Data Efficiency and Scalability Analysis.
4.2 Data Efficiency and Scalability
To analyze data efficiency, we pre-trained ARCE on increasingly larger subsets (25%, 50%, 75%, and 100%) of the Cote corpus. As illustrated in Figure 3, model performance scales almost linearly with the volume of generated data. The Macro-F1 score steadily increases from 74.23% using 25% of the corpus to 77.20% with the full dataset. Notably, the performance curve shows no signs of saturation, suggesting that further gains are achievable with an even larger Cote corpus. This sustained growth trajectory indicates that the model’s strength derives from the cumulative effect of a large and diverse set of elucidations, rather than from a few ”golden” samples, highlighting the robust scalability of our approach.
5 Conclusion
In this study, we introduced ARCE, a novel approach that enhances a RoBERTa model with LLM-generated knowledge for NER in the specialized AEC domain. Our method establishes a new state-of-the-art and, more importantly, reveals a ”less is more” principle for knowledge generation: a large corpus of simple, direct explanations is more effective for domain adaptation than complex, role-based rationales. While validated on a single dataset, future work will proceed in two directions: (1) evaluating ARCE across broader domains to test its generalizability, and (2) integrating the module into an end-to-end ARC system to measure the downstream impact on automated compliance checking.